OpenStudy (anonymous):

Simplify the product. 5b (6b-10)

OpenStudy (anonymous):

\[= 5b \times 6b - 5b \times 10 = ??\]

OpenStudy (anonymous):

\[\huge 30b^2-50b\]

OpenStudy (cwrw238):

you apply the Distributive law: multiply 6b and then -10 by the term outside ( 5b)